# Faster Language Models with Better Multi-Token Prediction Using Tensor Decomposition

### v1

@@ 3.3 Inference time benchmarking
Figure 5: Speculative decoding performance for trained head of the PyCode transformer model with different CP-rank values.
Table 2: Average number of accepted draft tokens for the PyCode model.
Rank Loss Average Draft Tokens Accepted
1 2.07 1.52
3 1.88 1.64
5 1.80 1.65
@@
To determine the impact of modified head on the inference time of bigger language models we benchmarked the time of one forward pass of our approach on large-scale models with 3 billion and 8 billion parameters. As reported in Table 3, the inference overhead for integrating the proposed multi-head layer remains minimal, even as the rank increases. Note that for the value of CP-rank of $5$ we observe a significant time increase for head execution then the Llama model is considerd which is probably caused by its huge vocabulary size. However, for moderate-sized networks inference time remains limited and increases only slightly with increasing CP-rank. The obtained results correspond to the theoretical algorithmic complexity of our new layer. During inference computational complexity of barebone grows linearly (given KV caches), but computational complexity of rank-$r$ head is always the same. Our measurements were made with seq length varying from 1024 to 4096, but for many practical applications sequence length is bigger, which further justifies usage of rank-$r$ head in the case of models with a large context window.
@@ 4 Related work
Training large language models (LLMs) to predict multiple tokens all at once and in parallel can drive these models toward better sample efficiency. Various approaches for multi-token predictions have been proposed recently. In Stern et al. (2018) several feed-forward decoder layers from the last encoder state are added for prediction of the next several tokens simultaneously, and in Miao et al. (2024) this idea was further improved within the framework of the so-called Medusa heads that use tree attention mechanism. In a number of works Song et al. (2021); Santilli et al. (2023); Fu et al. (2024) it is proposed to generate multiple draft tokens in parallel on the basis of the Jacobi iteration methods, i.e., via solving a non-linear system of equations while auto-regressive decoding in LLM. In the work Bhendawade et al. (2024) multiple tokens are predicted by adding streaming embeddings initialized from upper layers, with the token tree reduced by early exiting.
@@
Thus, this direction of research is actively developing today, however, the approaches outlined above have several limitations, including the need for significant changes in the original architecture of the model and limited speedup. Therefore, of particular interest is the recent work Gloeckle et al. (2024), where it was proposed to approximate the joint probability of the next several tokens using multiple heads but assuming conditional independence given the previous context. As we have already noted above, this approach remains a crude approximation that may limit token acceptance rates in the speculative decoding approach due to its disregard for token interdependencies. To improve upon this, we considered in this work a more accurate approximation of the joint distribution in the form of the CP-decomposition.
@@
To effectively implement the proposed scheme, we paid attention to connection of the used weighted CP-decomposition with the Mixture of Experts (MoE) technique. MoE is a widespread approach to enhance capabilities of LLMs with the most popular one being Sparse-Gated MoE introduced in Shazeer et al. (2017). MoE implementations can be either sparse or dense with sparse version being more popular, but there are many usages of both options, as in Dou et al. (2023) and Pan et al. (2024). While many parameters of MoE approach can be tweaked (Cai et al., 2024), the most common option is using MoE inside a transformer block, as in Zhou et al. (2022). We also note that MoE usage is not limited to LLMs and, for example, in Oldfield et al. (2024) it is applied to computer vision model.
@@
In this work, as an application of the proposed model for multi-token prediction, we consider its use as part of the speculative decoding scheme, which was proposed in Leviathan et al. (2023) and nowadays has become a common technique in the domain of inference acceleration. While initial framework solves the problem of inference optimization of a model given a faster draft model, there are different methods to obtain this draft model. Early works proposed blockwise decoding as in Stern et al. (2018). This line of work is similar to ours, as the model, used for speculative decoding, is exactly the same, as base model. Later more techniques for self-speculative decoding were developed, namely in Elhoushi et al. (2024) it is proposed to use only particular layers of the base model to obtain draft model and in Hooper et al. (2023) the base model consists of cycles, which also allows to skip layers to obtain a draft model. Self speculative decoding and multi-token prediction naturally go well with each other. This combination may require modification in model architecture as in Bhendawade et al. (2024), but it is possible to modify only heads as in Gloeckle et al. (2024) to enable faster application of the approach to existing LLMs, and we use such approach in our work.

### v2

@@ 4.3 Inference time benchmarking
Figure 6: Speculative decoding performance for fune-tuned head of 2B transformer model on fineweb-edu dataset with different CP-rank values.
Table 3: Average number of accepted draft tokens and benchmark time for fune-tuned model.
Rank Execution time (s) Avg. Draft Tokens
1 39.3 2.10
2 36.9 2.32
4 36.2 2.38
@@
To determine the impact of modified head on the inference time of bigger language models we benchmarked the time of one forward pass of our approach on large-scale models with 3 billion and 8 billion parameters. As reported in Table 4, the inference overhead for integrating the proposed multi-head layer remains minimal, even as the rank increases. Note that for a CP-rank value of 5, we observe a significant increase in head execution time when considering the Llama model, likely due to its large vocabulary size. However, for moderately sized networks, inference time remains manageable and increases only slightly with higher CP-rank values. The obtained results align with the theoretical algorithmic complexity of our new layer. During inference, the computational complexity of the barebone model grows linearly (given KV caches), whereas the computational complexity of a rank-$r$ head remains constant. Our measurements were conducted with sequence lengths varying from 1024 to 4096, but in many practical applications, sequence lengths are even longer, further justifying the use of rank-$r$ heads in models with large context windows.
@@ 5 Related work
Training large language models (LLMs) to predict multiple tokens all at once and in parallel can drive these models toward better sample efficiency. Various approaches for multi-token predictions have been proposed recently. In (Stern et al., 2018) several feed-forward decoder layers from the last encoder state are added for prediction of the next several tokens simultaneously, and in (Miao et al., 2024) this idea was further improved within the framework of the so-called Medusa heads that use tree attention mechanism. In a number of works (Song et al., 2021; Santilli et al., 2023; Fu et al., 2024) it is proposed to generate multiple draft tokens in parallel on the basis of the Jacobi iteration methods, i.e., via solving a non-linear system of equations while auto-regressive decoding in LLM. In the work (Bhendawade et al., 2024) multiple tokens are predicted by adding streaming embeddings initialized from upper layers, with the token tree reduced by early exiting.
@@
Thus, this direction of research is actively developing today, however, the approaches outlined above have several limitations, including the need for significant changes in the original architecture of the model and limited speedup. Therefore, of particular interest is the recent work (Gloeckle et al., 2024), where it was proposed to approximate the joint probability of the next several tokens using multiple heads but assuming conditional independence given the previous context. As we have already noted above, this approach remains a crude approximation that may limit token acceptance rates in the speculative decoding approach due to its disregard for token interdependencies. To improve upon this, we considered in this work a more accurate approximation of the joint distribution in the form of the CP-decomposition.
@@
To effectively implement the proposed scheme, we paid attention to connection of the used weighted CP-decomposition with the Mixture of Experts (MoE) technique. MoE is a widespread approach to enhance capabilities of LLMs with the most popular one being Sparse-Gated MoE introduced in (Shazeer et al., 2017). MoE implementations can be either sparse or dense with sparse version being more popular, but there are many usages of both options, as in (Dou et al., 2023) and (Pan et al., 2024). While many parameters of MoE approach can be tweaked (Cai et al., 2024), the most common option is using MoE inside a transformer block, as in (Zhou et al., 2022).
@@
In this work, as an application of the proposed model for multi-token prediction, we consider its use as part of the speculative decoding scheme, which was proposed in (Leviathan et al., 2023) and nowadays has become a common technique in the domain of inference acceleration. Self speculative decoding and multi-token prediction naturally go well with each other. This combination may require modification in model architecture as in (Bhendawade et al., 2024), but it is possible to modify only heads as in (Gloeckle et al., 2024; Li et al., 2024) to enable faster application of the approach to existing LLMs, and we use such approach.